Sally is Bob’s younger sister, and she is 7 years younger than Bob. If you multiply Bob’s age by three and subtract Sally’s age, you get their mother’s age, which is 37 years. How old is Sally?

Answer by stanbon(75887)

Sally is Bob’s younger sister, and she is 7 years younger than Bob. If you multiply Bob’s age by three and subtract Sally’s age, you get their mother’s age, which is 37 years. How old is Sally?
------
Equations:
s = b - 7
3b - 2s = 37
------
Substitute for "b" and solve for "s":
3(s+7) - 2s = 37
s + 21 = 37
s = 16 yrs. (Sally's age now)
=====================